Determine the long-run total, average and marginal cost functions of the firms having the
production functions and facing the input prices given below:
a. ?=f(K,L)=√K+2√L,PK=1,PL=2
b. Q=f(K,L)=K+L,PK=2,PL=1
c. Q=f(K,L)=K1/2L1/2,PK=2,PL=2
Answer #1

fck L) = VK +2 VE 0410 Lok%2+242 a Q= f (K; L) = VEL Now, Q- FCK,L) - KV2+2, s Now, marginal product of capital is.fi duct oNow long run total cost brition mit a renta will be P Lt Pk kietoa LEDEL ESTAD 021Q7 AUD We knool2O3p E a Te = 2Q z bogate peHuse, Q= Lto Just 107 -) Q-L and Totul cost siste puctions tipi Ko ahun ku ini sering PLLT Pk K TC= tot Aidha - Q + (axo)*:!Now, Q=K²L/ substituting K= L we get, a=2²2² As K= L, Huce K=Q Hace Long run total cost for ction LTC = PLL + Pk ka =2Q + QQ.